OpenStudy (gtxmuqsit):

A currebt of 8 amperes is flowing in a wire. How many electrons pass a given point int he wire in 1 second, if the charged on an electron is 1.6x10 to the power -19 coloumb? a)3.5x10^19 electrons b)5x10^19 electrons c)7.25x10^19 electrons d)2.75x10^15 electrons e)3.25x10^17 electrons

OpenStudy (gtxmuqsit):

@AllTehMaffs @Mashy @mathslover @Lessis @sarah786

OpenStudy (anonymous):

An amp is a Coulomb per second, so you can set up a conversion \[\Bigg( \frac{8A}{} \Bigg) \Bigg( \frac{1 C}{1 s}\Bigg) \Bigg( \frac{1 e^-}{1.6\times10^{-19}C}\Bigg) = \frac{\# \ e^{-}}{s} \]

OpenStudy (anonymous):

\[I = \frac{ Q }{ \Delta t }\] \[I \Delta t = Q\] \[\frac{ Total charge }{ charge per electron} = number of electrons\]

OpenStudy (gtxmuqsit):

@AllTehMaffs @Lessis thanks alot :P

OpenStudy (anonymous):

^_^